OpenStudy (anonymous):

If (x + y) dx + (x – y) dy = 0, then (a) x^2 + 2xy - y^2=C (b) x^2 - 2xy - y^2=C (c) x^2 + 2xy + y^2=C (d) x^2 - 2xy + y^2=C

OpenStudy (loser66):

I solved and got the answer not one of your choices, mine is \(x^2+4xy -y^2 =c\) while your a has just 2 xy. So, don't know

OpenStudy (anonymous):

\[(x+y)+(x-y)y'=0\] If \(M(x,y)=x+y\) and \(N(x,y)=x-y\), check if the equation is exact: \[\frac{\partial M}{\partial y}=1~~~~~~~~~~\frac{\partial N}{\partial x}=1\] The partial derivatives are equivalent, so you have an exact equation. So now you have \(\Psi_x=M(x,y)=x+y.\) \[\Psi_x=x+y\\ \int\Psi_x~dx=\int(x+y)~dx\\ \Psi=\frac{1}{2}x^2+xy+h(y)\] Differentiate \(\Psi\) w.r.t to y: \[\Psi_y=N(x,y)\\ x+h'(y)=x-y\\ h'(y)=-y\\ h(y)=-\frac{1}{2}y^2+C\] So the solution is \(\Psi=C\): \[\frac{1}{2}x^2+xy-\frac{1}{2}y^2=C\\ x^2+2xy-y^2=C\]